OpenStudy (anonymous):

What is the y-intercept for this linear equation? 1/2x-2/3y=-6

OpenStudy (precal):

solve for y

OpenStudy (precal):



OpenStudy (luigi0210):

plug in 0 for x and solve

OpenStudy (luigi0210):

and don't lose your negative!

OpenStudy (precal):

yes that probably will be much faster

OpenStudy (anonymous):

so its 1/2(0)-2/3y=-6 and do i divieded by what

OpenStudy (luigi0210):

well you don't exactly divide when dealing with fractions.. you multiply by it's reciprocal

OpenStudy (anonymous):

how is that

OpenStudy (luigi0210):

\[-\frac{ 2 }{ 3 }y=-6\] Now the reciprocal would be -3/2 to cancel out the negative and fraction

OpenStudy (luigi0210):

\[(-\frac{ 3 }{ 2 })(-\frac{ 2 }{ 3 }y)=-6(-\frac{ 3 }{ 2 })\]

OpenStudy (anonymous):

the what do i do when i cancle the fractions

OpenStudy (luigi0210):

It becomes \[y=\frac{ 18 }{ 2 }=9\]

OpenStudy (anonymous):

ok thanks

OpenStudy (luigi0210):

Yup!

OpenStudy (precal):